OpenStudy (anonymous):

Mike started a savings account by depositing $9. Each month he deposits more money than the month before. At the end of 41 months, he has saved $9,989.00. How much more does he deposit each month? A.$11.00 B.$11.50 C.$12.00 D.$12.50

OpenStudy (texaschic101):

I am having trouble understanding this...if each month he deposits more money then the month before, then he is not depositing the same amount of money each month...right ?

OpenStudy (anonymous):

Yes because its an Arithmetic Series

OpenStudy (tkhunny):

There is insufficient information in the problem statement to state that it is an arithmetic series. Is it really? It might say that, I suppose.

OpenStudy (anonymous):

Well the topic for this lesson is "08.02 Arithmetic Series" and i just copied and pasted the question

OpenStudy (tkhunny):

Fair enough. It should be written more clearly. Any way... How do you propose we proceed with this? Have you a formula for a sum of 41 payments in an arithmetic series or shall we build it?

OpenStudy (texaschic101):

Are your answer choices supposed to be the common differences ?

OpenStudy (anonymous):

The formula is to find the sum is Sn=n/2( a1 + an)

OpenStudy (anonymous):

and I'm guessing the answer choices is the amount of money he deposits each month

OpenStudy (anonymous):

there it is the formula

OpenStudy (anonymous):

No i am pretty sure you're right. I just don't know how we're suppose to figure out how much he put in each month

OpenStudy (anonymous):

Lets up so

OpenStudy (anonymous):

hope**

OpenStudy (texaschic101):

the only way I am getting 469 as the 41st term is if I use 11.50 as the common difference But when I add the sums of the 41st terms, I am ending up with 9799 instead of 9989.

OpenStudy (texaschic101):

and when I use 12 as the common difference, I am ending up with over 10,000 as the sum of 41 terms.

OpenStudy (anonymous):

I'll just guess its only one question But thank you guys so much for trying to solve it

OpenStudy (texaschic101):

I am sorry that I could not help :(

OpenStudy (tkhunny):

Let N be the common difference in the payments. Add them up! 9 9+N 9+2N 9+3N ... 9+40N 9+41N That's 42 deposits. 9(42) + N(1 + 2 + 3 + ... + 40 + 41) = 378 + N((41*42)/2) = 378 + N(41*21) = 378 + 861N Finally: 378 + 861N = 9989 and N = 11.1626 I know, you're thinking, "Why did we go all the way to 42 payments?" Well, suppose we don't and we cut off at the slightly more motivated 41 payments... 369 + 820N = 9989 and N = 11.7317 Where does that leave us? I show: 41 Payments 11.00 ==> 9389 11.50 ==> 9799 12.00 ==> 10209 42 Payments 11.00 ==> 9849 11.50 ==> 10279.50 So, $9989.00 ? I'm not seeing it.

OpenStudy (anonymous):

I put 11.50$ as the answer and it was incorrect :/

OpenStudy (anonymous):

Its no problem! thank you everyone

OpenStudy (texaschic101):

hmm....I got 11.50 as the closest answer, and when I used 12 as the common difference, it was over 10,000.

OpenStudy (tkhunny):

Except that makes no sense. We are adding whole dollars or half dollars. It is not acceptable for ANY deposit to end with 0.265. 11.50 is no good. 9 + 40*11.50 = 469 This is not a good question. Something seriously wrong with it.

OpenStudy (anonymous):

I took the test and the correct answer is $11.00